OpenStudy (anonymous):

In December, A, B, C and D saved a total of $1267. In January, each of them had an equal amound of savings when A's savings was doubled, B's was increased by $18 and D's was decreased by $25. If C's remained the same, what was B's savings in January?

OpenStudy (anonymous):

*amount

OpenStudy (anonymous):

A+B+C+D=1267 A*2=B+18 => A=(B+18)/2 B+18=C B+18=D-25 => D=B+18+25=B+43 put eveyrhting into first (B+18)/2+B+B+18+B+43=1267 B=342

OpenStudy (anonymous):

\[\frac{9}{2}B + $70 = $ 1267\] \[\frac{9}{2}B = $(1267-70) = $ 1197\] B = $ 262

OpenStudy (anonymous):

i feel so so in jan B's saving is $ 280

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

hmm i calculated B ammounf of money in decemeber so in january it would be 342+18=360